OpenStudy (yrivers36):

Find the domain and write in interval notation sqrt of (10-2x)

OpenStudy (anonymous):

since you can't square a negative number, you must first solve for x

OpenStudy (anonymous):

10-2x=0, 10=2x, x=5

OpenStudy (savvy):

\[x \epsilon[5,\infty)\]

OpenStudy (savvy):

since \[10 - 2x \ge 0\]

OpenStudy (yrivers36):

ok thank you. I thought so but was not sure.

OpenStudy (savvy):

welcome...!!!!

OpenStudy (yrivers36):

would it be x is less than or equal to 5